# The birth mass function of neutron stars

### v1

@@ B.2 A phenomenological model
Here we estimate the accreted masses of recycled neutron stars using a phenomenological relation between $\Delta m$ and the initial spin periods $P_{\rm{ini}}$. As in the analytic model, for group $b$ neutron stars, we assume that the observed spin periods are their initial spin periods. For group $a$ pulsars, the initial spin period is determined as follows. We assume that the true age of a recycled pulsar is uniformly drawn from $[0,\tau_{\rm{max}}]$. For a given true age, we evolve the pulsar back in time following the spin-down equation $\dot{P}\sim P^{2-n}$ to find its $P_{\rm{ini}}$. We generate $10^{4}$ random ages uniformly drawn in $[0,\tau_{\rm{max}}]$, and obtain the distribution of initial spin periods $P_{\rm{ini}}$. For a given $P_{\rm{ini}}$, $\Delta m$ is drawn from a Gaussian distribution with a mean of $\mu_{M}/\mathrm{M}_{\odot}=0.3P_{\rm ms}^{-1/2}$ (with $P_{\rm ms}$ being the initial spin period measured in millisecond) and a standard deviation of $0.2\mu_{M}$. The mean in this recipe is motivated by numerical simulations that take into account the neutron star spin-down due to propeller effects during the recycling process[28]; see blue dots Extended Data Figure 2. The $20\%$ dispersion is the typical spread found in the analytical prescription.
@@
Extended Data Figure 2 demonstrates our prescriptions for the accreted mass-spin period correlation. The green and orange solid lines represent $\Delta m_{\rm min}/\mathrm{M}_{\odot}=0.22P_{\rm ms}^{-4/3}$ and $\mu_{M}/\mathrm{M}_{\odot}=0.3P_{\rm ms}^{-1/2}$, respectively. The grey and orange shaded bands encompass 90% credible region of the analytical and phenomenological model, respectively. The red dashed line is $\Delta m/\mathrm{M}_{\odot}=0.2P_{\rm ms}^{-1}$. Results from numerical simulations of the recycling process in ref.[28] are marked as blue dots; note that their calculations were performed for a fiducial value of initial magnetic field strength and neutron star moment of inertia. Our models captures the overall trend and scatter of numerical results well. The difference between grey and orange bands at $P\gtrsim 10\,\mathrm{ms}$ appears to be exaggerated due to the logrithmic scale of the vertical coordinate of Extended Data Figure 2.
@@
B.3 A case study of PSR J1614--2230
PSR J1614--2230, with an observed spin period of $3.15\,\mathrm{ms}$, was the first binary pulsar having a measured neutron star mass around $2\,\mathrm{\rm M_{\odot}}$. The pulsar mass was originally measured as $1.97\pm 0.04\,\mathrm{\rm M_{\odot}}$, and the neutron star is found in an 8.7-day orbit with a $0.5\,\mathrm{\rm M_{\odot}}$ white dwarf companion[10]. By performing detailed binary evolution calculations, Tauris et al.[85] suggested that the most likely progenitor system of PSR J1614--2230 is a $1.7\pm 0.15\,\mathrm{\rm M_{\odot}}$ neutron star going through “Case A” Roche lobe overflow (and thus accreting mass) from a $\sim 4.5\,\mathrm{\rm M_{\odot}}$ donor star. An independent binary evolution study found a minimum initial (birth) mass of $1.6\pm 0.1\,\mathrm{\rm M_{\odot}}$[86].
The mass measurement of PSR J1614--2230 was updated in 2018 to a slightly lower value of $1.908\pm 0.016\,\mathrm{\rm M_{\odot}}$ with extended observations[31]. From our accreted-mass modelling, we find its birth mass is around $1.7\,\mathrm{\rm M_{\odot}}$, in excellent agreement with binary evolution calculations. Extended Data Figure 3 shows the measured and inferred birth mass distributions of PSR J1614--2230. The blue and orange curves indicate the birth masses estimated from our analytical and phenomenological modelling, respectively.
@@ C Bayesian inference
In this section, we briefly describe the method of Bayesian inference used to determine the birth mass function of neutron stars; a more general introduction can be found in refs.[87, 88].
The data being analyzed are posterior distributions for 90 neutron stars. The original data set, denoted as OBS, consists of Gaussian mass posteriors used to approximate measurements reported in the literature (as summarized in Extended Data Tables 1 and 2) or posterior samples if available. For 53 recycled pulsars in our compilation, we also apply probabilistic accreted mass corrections to obtain posteriors for their birth masses. The data set obtained with the analytical (phenomenological) prescription described in Section B.1 is denoted as ANA (PHE). A data set that includes only those measurements of slow (nonrecycled) neutron stars is denoted as SLOW. For comparison, the data set that includes only observed masses of 53 recycled pulsars is denoted as REC. The $\rm{ANA}_{m_{r}}$ ($\rm{PHE}_{m_{r}}$) data set includes only birth masses of 53 recycled pulsars after applying the analytical (phenomenological) accreted-mass correction. The ANAradio data set is a subset of ANA where we only include mass measurements obtained by the radio pulsar timing method (i.e., all in Extended Data Table 1, plus PSR J1903+0327).
Let $p(m^{i}|d^{i})$ be the mass posterior distribution of the $i$-th neutron star, given some observations $d^{i}$. Our goal is to find a model that can best describe the collection of mass posteriors for $N$ sources. We introduce conditional prior for neutron star mass $\pi(m^{i}|\Lambda)$ for a given model $M$ that includes hyperparameters $\Lambda$. The hyperposterior distribution of $\Lambda$ can be written as
@@
where $\{d\}$ denotes the original observations that provide individual neutron star mass measurements, $\pi(\Lambda|M)$ is the prior distribution for $\Lambda$, and the term ${\cal L}(d^{i}|m^{i})$ is the likelihood function of the data given $m^{i}$, which is related to the mass posterior via
@@
where $\pi_{0}(m^{i})$ is the initial prior used to derive the mass posterior (assumed to be flat here), ${\cal Z}_{0}^{i}$ is the initial evidence which is a factor that will ultimately cancel via either normalization of the hyperposterior or the construction of a Bayesian evidence ratio.
@@
In practice, we can replace the integral over ${\rm d}m^{i}$ with a summation over evaluating the conditional prior at posterior samples of $m^{i}$, so that the hyperevidence is given by
@@
where $n_{i}$ is the number of mass posterior samples for the $i$-th neutron star.
In order to find the model that best fits the birth mass function of neutron stars, we compute the Bayes factor between two models ($M_{1}$ and $M_{2}$), defined as follows
@@
We follow the interpretation of Bayes factors outlined by ref.[89], and choose a threshold for BF of 150, corresponding to a natural logarithmic BF of 5, as required for confident model selection. Note that the Bayes factor is equal to the odds, defined as the ratio of posterior probability between two models, assuming that their prior probabilities are equal.
The likelihood function adopted above is applicable to the case where individual neutron star mass measurements are independent. For Galactic double neutron star systems, the mass measurements are correlated because of the significant constraint on the total mass of two neutron stars, as demonstrated in fig. 1 in ref.[90]. In this case, the joint posteriors of two masses should be used following the formalism outlined in section 3.1 of ref.[90]. To illustrate this point explicitly, the total likelihood that is hidden in Equation (8) is
@@
which should be modified as follows
@@
Here 90 neutron stars are divided into two groups: the first group include 64 neutron stars whose mass posteriors are truly independent; the second group contains 13 pairs of neutron stars (labeled as $m_{1}^{j}$ and $m_{2}^{j}$) that are in double neutron star systems. For the second group, the individual likelihood function ${\cal L}(d^{j}|m_{1}^{j},m_{2}^{j})$ is proportional to the joint posterior $p(m_{1}^{j},m_{2}^{j}|d^{j})$.
In our analysis, we make use of the software package BILBY[91] for parameter estimation and evidence calculation, and the ${\tt{gwpopulation}}$[92] package for population modelling.
@@ D Parametric models for the neutron star mass function
Our base model is a uniform distribution
@@
where $m_{\min}$ and $m_{\max}$ are the maximum and minimum neutron star mass, respectively. For the uniform model and all other models where applicable, we adopt uniform priors in the range of $[0.9,2.9]\mathrm{M}_{\odot}$ and $[1.9,2.9]\mathrm{M}_{\odot}$ for $m_{\min}$ and $m_{\max}$, respectively. The prior edges are chosen to be the same as in ref.[16], to facilitate comparisons.
@@
Here the power-law index takes a uniform prior between $-5$ and 25. The power-law distribution can be modified to allow a smooth turn-on over a range of $\delta m$, which takes a uniform prior between 0.01 and $1\mathrm{M}_{\odot}$ in our analysis. Such a turn-on power-law model (TOP) is defined as
@@
where $\mathcal{H}$ is the Heaviside step function, $S$ denotes a smoothing function which is designed to rise from 0 at $m_{\min}$ to 1 at $(m_{\min}+\delta m)$. In the interval of $m_{\min}<m<m_{\min}+\delta m$, the function $S$ can be expressed by
@@
Note that in plots showing the posterior distributions of TOP parameters, we show the parameter $m_{\rm peak}$ which defines the peak location of the distribution.
@@
where $\Gamma(\alpha)$ is the gamma function. We adopt uniform priors for $\alpha$ and $\theta$ in the range of $(0,80)$ and $(0.01,0.1)$, respectively.
The skewed Student’s t (SST) distribution can be parameterized as[93, 94]
@@
Here $\mu$ and $\sigma$ are the mean and standard deviation of the Gaussian distribution, respectively. The parameters $\nu$ and $\tau$ are used to characterize the degree of skewness. In Equation (23), $\mu_{0}=$ $\mu-\sigma\kappa/s$, $\sigma_{0}=\sigma/s$, $z=\left(m-\mu_{0}\right)/\sigma_{0}$, and $c=2\nu\left[\left(1+\nu^{2}\right)B(1/2,\tau/2)\tau^{1/2}\right]^{-1}$ with $\kappa$ and $s$ given by $\kappa=2\tau^{1/2}\left(\nu-\nu^{-1}\right)[(\tau-1)B(1/2,\tau/2)]^{-1}$ and $s=\left[\frac{\tau}{(\tau-2)}\left(\nu^{2}+\nu^{-2}-1\right)-\kappa^{2}\right]^{1/2}$, respectively; $B$ is the beta function. Priors for $\mu$ and $\sigma$ are the same as those of the Gaussian-family models, and we adopt uniform priors for $\nu$ and $\tau$ in the range of $(0,8)$ and $(2,20)$, respectively.
@@ F Semi-analytic supernova modeling
In this section we investigate the implications of the measured neutron star birth mass distribution for core-collapse supernova physics. Extended Data Figure 7a shows the distribution $\phi(M_{4})$ of the iron-silicon core mass $M_{4}$ (defined by a threshold value of $4\,\mathrm{k}_{\mathrm{B}}$ per nucleon in specific entropy), weighted by the initial mass function (IMF) for massive stars,[1, 2] for a set of 2912 single-star supernova progenitor models[21] with initial masses from $9.45\,\mathrm{\mathrm{M}_{\odot}}$ to $45\,\mathrm{\mathrm{M}_{\odot}}$. The parameter $M_{4}$ is a proxy for the mass coordinate for shock revival in neutrino-driven supernovae [22]. An additional factor $(M_{4}/M_{0})^{6.5}$ (with an arbitrary scaling factor $M_{0}$) is applied to highlight deviations from the fiducial $\phi(M_{4})\propto M_{4}^{-6.5}$ scaling where the power-law slope comes from the TOP model as determined for the birth mass function (see Figure 4).
@@
a)Refer to caption
b)Refer to caption
c)Refer to caption
Extended Data Figure 7: a) IMF-weighted distribution $\phi(M_{4})$ of the mass $M_{4}$ of the iron-silicon core for a set of 2912 single-star supernova progenitors of solar metallicity[21], with a compensation factor $(M_{4}/M_{0})^{6.5}$ to show deviations from the power-law of the measured neutron star mass distribution. b) Predicted distribution $\phi(M)$ of neutron star masses with the same compensation factor based on the semi-analytic supernova model of [21] with standard model parameters. c) Predicted distribution $\phi(M)$ of neutron star masses with modified supernova model parameters to allow for more accretion after shock revival.
a)Refer to caption b)Refer to caption
c)Refer to caption d)Refer to caption
Extended Data Figure 8: a) Predicted IMF-weighted neutron star mass distribution $\phi(M)$ using the semi-analytic supernova model of [21] with standard model parameters. b) Distribution of the amount $\Delta M$ of net accretion after shock revival. c) and d) Predicted IMF-weighted distribution of neutron star masses and accretion after shock revival for the modified supernova model with enhanced accretion during the explosion phase.
Extended Data Figure 9: a) Explosion energy, b) black-hole masses for non-exploding cases and fallback supernovae, c) neutron star masses, d) nickel masses predicted by the semi-analytic supernova model with adjusted parameters for enhanced accretion to better match the observed power-law distribution of neutron star masses. Red dots in panels a) and d) denote explosions with black hole formation by fallback. Explosion energies and nickel masses span the typical range observed for Type IIP supernovae.

### v2

@@ B.2 A phenomenological model
Here we estimate the accreted masses of recycled neutron stars using a phenomenological relation between $\Delta m$ and the initial spin periods $P_{\rm{ini}}$. As in the analytic model, for group $b$ neutron stars, we assumed that the observed spin periods are their initial spin periods. For group $a$ pulsars, the initial spin period was determined as follows. We assumed that the true age of a recycled pulsar is uniformly drawn from $[0,\tau_{\rm{max}}]$. For a given true age, we evolvd the pulsar back in time following the spin-down equation $\dot{P}\propto P^{2-n}$ to find its $P_{\rm{ini}}$. We generated $10^{4}$ random ages uniformly drawn from $[0,\tau_{\rm{max}}]$ and obtained the distribution of initial spin periods $P_{\rm{ini}}$. For a given $P_{\rm{ini}}$, $\Delta m$ is drawn from a Gaussian distribution with a mean of $\mu_{M}/\mathrm{M}_{\odot}=0.3P_{\rm ms}^{-1/2}$ (with $P_{\rm ms}$ being the initial spin period measured in milliseconds) and a standard deviation of $0.2\mu_{M}$. The mean in this recipe is motivated by numerical simulations that take into account the neutron star spin-down due to propeller effects during the recycling process[28]; see the blue dots Extended Data Figure 2. The $20\%$ dispersion is the typical spread found in the analytical prescription.
@@
Extended Data Figure 2 illustrates our prescriptions for the accreted mass-spin period correlation. The green and orange solid lines represent $\Delta m_{\rm min}/\mathrm{M}_{\odot}=0.22P_{\rm ms}^{-4/3}$ and $\mu_{M}/\mathrm{M}_{\odot}=0.3P_{\rm ms}^{-1/2}$, respectively. The grey- and orange-shaded bands encompass the 90% credible regions of the analytical and phenomenological models, respectively. The red dashed line is $\Delta m/\mathrm{M}_{\odot}=0.2P_{\rm ms}^{-1}$. Results from numerical simulations of the recycling process in ref.[28] are marked as blue dots. Note that their calculations were performed for a fiducial value of the initial magnetic field strength and neutron star moment of inertia. Our models captures the overall trend and scatter of the numerical results well. The difference between grey and orange bands for $P\gtrsim 10\,\mathrm{ms}$ is exaggerated due to the logarithmic scale of the vertical coordinate in Extended Data Figure 2.
@@
B.3 A case study of PSR J1614--2230.
PSR J1614--2230, which has an observed spin period of $3.15\,\mathrm{ms}$, was the first binary pulsar found to have a measured neutron-star mass around $2\,\mathrm{\rm M_{\odot}}$. The pulsar mass was originally measured as $1.97\pm 0.04\,\mathrm{\rm M_{\odot}}$, and the neutron star is in an 8.7-day orbit with a $0.5\,\mathrm{\rm M_{\odot}}$ white dwarf companion[10]. By performing detailed binary evolution calculations, Tauris et al.[85] suggested that the most probable progenitor system of PSR J1614--2230 is a $1.7\pm 0.15\,\mathrm{\rm M_{\odot}}$ neutron star going through “case A” Roche lobe overflow (and, thus, accreting mass) from an $\sim 4.5\,\mathrm{\rm M_{\odot}}$ donor star. An independent binary evolution study found a minimum initial (birth) mass of $1.6\pm 0.1\,\mathrm{\rm M_{\odot}}$ (ref.[86]).
The mass measurement of PSR J1614--2230 was updated in 2018 to a slightly lower value of $1.908\pm 0.016\,\mathrm{\rm M_{\odot}}$ with extended observations[31]. From our accreted-mass modelling, we found that its birth mass was around $1.7\,\mathrm{\rm M_{\odot}}$, in excellent agreement with binary evolution calculations. Extended Data Figure 3 shows the measured and inferred birth-mass distributions of PSR J1614--2230. The blue and orange curves indicate the birth masses estimated from our analytical and phenomenological modelling, respectively.
@@ C Bayesian inference
In this section, we briefly describe the method of Bayesian inference used to determine the birth-mass function of neutron stars; a more general introduction can be found in refs.[87, 88].
The data analyzed are posterior distributions for 90 neutron stars. The original dataset, denoted as OBS, consists of Gaussian mass posteriors used to approximate measurements reported in the literature (as summarized in Extended Data Tables 1 and 2) or posterior samples, if available. For 53 of the recycled pulsars in our compilation, we also applied probabilistic accreted mass corrections to obtain posteriors for their birth masses. The dataset obtained with the analytical (phenomenological) prescription described in Section B.1 is denoted as ANA (PHE). A dataset that includes only those measurements of slow (non-recycled) neutron stars is denoted as SLOW. For comparison, the dataset that includes only observed masses of 53 recycled pulsars is denoted as REC. The $\rm{ANA}_{m_{r}}$ ($\rm{PHE}_{m_{r}}$) dataset includes the birth masses of 53 recycled pulsars only after applying the analytical (phenomenological) accreted-mass correction. The ANAradio dataset is a subset of ANA that includes only mass measurements obtained by the radio pulsar timing method (i.e., all in Extended Data Table 1, plus PSR J1903+0327).
Let $p(m^{i}|d^{i})$ be the mass posterior distribution of the $i$th neutron star, given some observations $d^{i}$. Our goal is to find a model that can best describe the collection of mass posteriors for $N$ sources. We introduce a conditional prior for neutron-star mass $\pi(m^{i}|\Lambda)$ for a given model $M$ that includes hyperparameters $\Lambda$. The hyperposterior distribution of $\Lambda$ can be written as
@@
where $\{d\}$ denotes the original observations that provide individual neutron-star mass measurements, $\pi(\Lambda|M)$ is the prior distribution for $\Lambda$ and the term ${\cal L}(d^{i}|m^{i})$ is the likelihood function of the data given $m^{i}$, which is related to the mass posterior through
@@
where $\pi_{0}(m^{i})$ is the initial prior used to derive the mass posterior (assumed to be flat here) and ${\cal Z}_{0}^{i}$ is the initial evidence, which is a factor that will ultimately cancel out through either normalization of the hyperposterior or the construction of a Bayesian evidence ratio.
@@
In practice, we can replace the integral over ${\rm d}m^{i}$ with a summation over the conditional prior at posterior samples of $m^{i}$, so that the hyperevidence is given by
@@
where $n_{i}$ is the number of mass posterior samples for the $i$th neutron star.
To find the model that best fits the birth-mass function of neutron stars, we compute the Bayes factor between two models ($M_{1}$ and $M_{2}$), defined as follows
@@
We followed the interpretation of Bayes factors outlined by ref.[89] and choose a threshold for BF of 150, corresponding to a natural logarithmic BF of 5, as required for confident model selection. Note that the Bayes factor is equal to the odds, which is defined as the ratio of the posterior probabilities between two models, assuming that their prior probabilities are equal.
The likelihood function adopted above is applicable when individual neutron-star mass measurements are independent. For Galactic double neutron star systems, the mass measurements are correlated because of the significant constraint on the total mass of two neutron stars, as demonstrated in figure 1 in ref.[90]. In this case, the joint posteriors of two masses should be used following the formalism outlined in Section 3.1 of ref.[90]. To illustrate this point explicitly, the total likelihood that is hidden in Equation (8) is
@@
which should be modified as follows:
@@
We divided the 90 neutron stars into two groups: the first group includes 64 neutron stars whose mass posteriors are truly independent; the second group contains 13 pairs of neutron stars (labelled as $m_{1}^{j}$ and $m_{2}^{j}$) that are in double neutron star systems. For the second group, the individual likelihood function ${\cal L}(d^{j}|m_{1}^{j},m_{2}^{j})$ is proportional to the joint posterior $p(m_{1}^{j},m_{2}^{j}|d^{j})$.
In our analysis, we used the software package BILBY (ref.[91]) for parameter estimation and evidence calculation and the ${\tt{gwpopulation}}$ package [92] for population modelling.
@@ D Parametric models for the neutron star mass function
Our base model is a uniform distribution:
@@
where $m_{\min}$ and $m_{\max}$ are the maximum and minimum neutron-star mass, respectively. For the uniform model and all other models where applicable, we adopted uniform priors in the ranges $[0.9,2.9]\mathrm{M}_{\odot}$ and $[1.9,2.9]\mathrm{M}_{\odot}$ for $m_{\min}$ and $m_{\max}$, respectively. The prior edges were chosen to be the same as in ref.[16] to facilitate comparisons.
@@
Here the power-law index takes a uniform prior between $-5$ and 25. The power-law distribution can be modified to allow a smooth turn-on over a range of $\delta m$, which takes a uniform prior between 0.01 and $1\mathrm{M}_{\odot}$ in our analysis. Such a TOP model is defined as
@@
where $\mathcal{H}$ is the Heaviside step function and $S$ denotes a smoothing function, which is designed to rise from 0 at $m_{\min}$ to 1 at $(m_{\min}+\delta m)$. In the interval of $m_{\min}<m<m_{\min}+\delta m$, the function $S$ can be expressed by
@@
Note that in plots of the posterior distributions of TOP parameters, we show the parameter $m_{\rm peak}$ which defines the peak location of the distribution.
@@
where $\Gamma(\alpha)$ is the gamma function, the $\alpha$ is the shape parameter and $\theta$ is the scalar parameter. We adopted uniform priors for $\alpha$ and $\theta$ in the range of $(0,80)$ and $(0.01,0.1)$, respectively.
The skewed Student’s $t$-distribution (SST) can be parameterized as[93, 94]
@@
Here $\mu$ and $\sigma$ are the mean and standard deviation of the Gaussian distribution, respectively. The parameters $\nu$ and $\tau$ are used to characterize the degree of skewness. In Equation (23), $\mu_{0}=$ $\mu-\sigma\kappa/s$, $\sigma_{0}=\sigma/s$, $z=\left(m-\mu_{0}\right)/\sigma_{0}$, and $c=2\nu\left[\left(1+\nu^{2}\right)B(1/2,\tau/2)\tau^{1/2}\right]^{-1}$ with $\kappa$ and $s$ given by $\kappa=2\tau^{1/2}\left(\nu-\nu^{-1}\right)[(\tau-1)B(1/2,\tau/2)]^{-1}$ and $s=\left[\frac{\tau}{(\tau-2)}\left(\nu^{2}+\nu^{-2}-1\right)-\kappa^{2}\right]^{1/2}$, respectively; $B$ is the beta function. Priors for $\mu$ and $\sigma$ were the same as those of the Gaussian-family models, and we adopted uniform priors for $\nu$ and $\tau$ in the ranges $(0,8)$ and $(2,20)$, respectively.
@@ F Semi-analytic supernova modeling
In this section we investigate the implications of the measured neutron star birth mass distribution for core-collapse supernova physics. Extended Data Figure 7a shows the distribution $\phi(M_{4})$ of the iron-silicon core mass $M_{4}$ (defined by a threshold value of $4\,\mathrm{k}_{\mathrm{B}}$ per nucleon in specific entropy), weighted by the initial mass function (IMF) for massive stars,[1, 2] for a set of 2912 single-star supernova progenitor models[21] with initial masses from $9.45\,\mathrm{\mathrm{M}_{\odot}}$ to $45\,\mathrm{\mathrm{M}_{\odot}}$. The parameter $M_{4}$ is a proxy for the mass coordinate for shock revival in neutrino-driven supernovae [22]. An additional factor $(M_{4}/M_{0})^{6.5}$ (with an arbitrary scaling factor $M_{0}$) is applied to highlight deviations from the fiducial $\phi(M_{4})\propto M_{4}^{-6.5}$ scaling where the power-law slope comes from the TOP model as determined for the birth mass function (see Figure 4 in the main text).
@@
aRefer to caption
bRefer to caption
cRefer to caption
Extended Data Figure 7: a, IMF-weighted distribution $\phi(M_{4})$ of the mass $M_{4}$ of the iron-silicon core for a set of 2912 single-star supernova progenitors of solar metallicity[21], with a compensation factor $(M_{4}/M_{0})^{6.5}$ to show deviations from the power-law of the measured neutron star mass distribution. b, Predicted distribution $\phi(M)$ of neutron star masses with the same compensation factor based on the semi-analytic supernova model of ref.[21] with standard model parameters. c, Predicted distribution $\phi(M)$ of neutron star masses with modified supernova model parameters to allow for more accretion after shock revival.
aRefer to caption bRefer to caption
cRefer to caption dRefer to caption
Extended Data Figure 8: a, Predicted IMF-weighted neutron star mass distribution $\phi(M)$ using the semi-analytic supernova model of ref.[21] with standard model parameters. b, Distribution of the amount $\Delta M$ of net accretion after shock revival. Panels c and d Predicted IMF-weighted distribution of neutron star masses and accretion after shock revival for the modified supernova model with enhanced accretion during the explosion phase.
Extended Data Figure 9: a) Explosion energy, b) black-hole masses for non-exploding cases and fallback supernovae, c) neutron star masses, d) nickel masses predicted by the semi-analytic supernova model with adjusted parameters for enhanced accretion to better match the observed power-law distribution of neutron star masses. Red dots in panels a and d denote explosions with black hole formation by fallback. Explosion energies and nickel masses span the typical range observed for Type IIP supernovae.